OpenStudy (anonymous):

trying to figure out phase shift for y = -1/2 sin (pi^2- Pi*x/2) ? I am completely stumped? I know what the graph looks like but I can't figure out whats' "C" or "B" values? (i.e. C/B)

OpenStudy (anonymous):

y = A sin[B (x+C) ] + D has the period of 2pi/B so your goal is to write the pi^2- Pi*x/2 in the form of B(x+C)

OpenStudy (anonymous):

so I would need to factor out a PI?

OpenStudy (jdoe0001):

or one can say \(\bf y=Asin(Bx+C)\qquad \textit{phase shift}\implies \cfrac{C}{B}\)

OpenStudy (anonymous):

So the "b" value is pi/2?

OpenStudy (anonymous):

but i'm still lost on how to put it into bx+c form?

OpenStudy (jdoe0001):

\(\bf y=-\frac{1}{2}sin\left(\pi^2-\frac{\pi}{2}x\right)\implies y=-\frac{1}{2}sin\left({\color{blue}{ -\frac{\pi}{2}}}x+{\color{blue}{ \pi^2}}\right) \\ \quad \\ \textit{phase shift}\implies \cfrac{{\color{blue}{ \pi^2}}}{{\color{blue}{ -\frac{\pi}{2}}}}\)

OpenStudy (anonymous):

so -2pi is the phase shift? I think i'm grasping straws:/

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

but when i calculated the period i got 4? and the graph doesn't look like that?...

OpenStudy (anonymous):

I think i missed step somewhere?

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

\(\bf y=-\frac{1}{2}sin\left(\pi^2-\frac{\pi}{2}x\right)\implies y=-\frac{1}{2}sin\left({\color{blue}{ -\frac{\pi}{2}}}x+{\color{blue}{ \pi^2}}\right) \\ \quad \\ period\implies \cfrac{2\pi}{{\color{blue}{ -\frac{\pi}{2}}}}\)

OpenStudy (jdoe0001):

actually you got the period correct :)

OpenStudy (anonymous):

I appreciate the help, I think maybe my graph is wrong, im not sure. I'm just going to try and keep re working. ty

OpenStudy (anonymous):

see, yes thats the exact same graph I had, but i'm confused as why the sine function is not intersecting at 0,0 even though its reflective across the x-axis, and the amplitude is 1/2, but yet it's like it starts it cycle at .2?